# ContentV: Efficient Training of Video Generation Models with Limited Compute

### v1

@@ 5.3 DiT Pre-Training
Stage 3 $125\times 432\times 768$ 2048 256 $5\times e^{-4}$ 30k
@@ 6.2 Reinforcement Learning from Human Feedback
RLHF for diffusion model aims to optimize a conditional distribution $p_{\theta}(x_{0}\mid c)$ such that the reward model $r(c,x_{0})$ defined on it is maximized, while regularizing the KL-divergence from a reference model $p_{\textrm{ref}}$. More specifically, RLHF optimizes a model $p_{\theta}$ to maximize the following objective:
$\max_{p_{\theta}}\mathbb{E}_{c\sim\mathcal{D}_{c},x_{0}\sim p_{\theta}(x_{0}\mid c)}[r(c,x_{0})]-\beta\mathbb{D}_{\mathrm{KL}}[p_{\theta}(x_{0}\mid c)\parallel p_{\mathrm{ref}}(x_{0}\mid c)]$ (2)
where the hyperparameter $\beta$ controls KL-regularization strength. Inspired by Wu et al. [50], supervises the final output of diffusion model and back-propagates through the iterative sampling process to the input noise. Instead of training at equal intervals, we randomly sample $k$ steps carrying gradient from the timesteps of infer. Considering that the model for video generation is large and the sequence is long while the memory of the AI accelerators is not large, we ignored the KL constraint for the sake of computational efficiency.
@@ 7.2 Asynchronous Encode Server
Figure 8: Asynchronous Encode Server. The training of the model and feature extraction are placed on different nodes, and the transmission of data and features is completed through asynchronous network requests.
Since the text-to-video diffusion model performs denoising in latent space, it requires both VAE and text encoder to extract latent features and text embeddings respectively. Although this inference stage involves no gradient computation, the T5-XXL [38] text encoder and VAE encoder still demand significant memory resources. In SD3.5L [12], feature extraction is performed offline, which inevitably consumes substantial storage space. To address these resource constraints, we propose a decoupled architecture that separates computing resources for inference and training. Specifically, the VAE and text encoder operate on dedicated inference clusters, continuously generating processed input data (including image/video latent featuers and corresponding text embeddings) for DiT training. The training cluster then retrieves these pre-computed features via network transmission and focuses exclusively on the training process.

### v2

@@ 5.3 DiT Pre-Training
Stage 3 $125\times 432\times 768$ 2048 256 $5\times e^{-5}$ 30k
@@ 6.2 Reinforcement Learning from Human Feedback
RLHF for diffusion model aims to optimize a conditional distribution $p_{\theta}(x_{1}\mid c)$ such that the reward model $r(c,x_{1})$ defined on it is maximized while regularizing the KL-divergence from a reference model $p_{\textrm{ref}}$. More specifically, RLHF optimizes a model $p_{\theta}$ to maximize the following objective:
$\max_{p_{\theta}}\mathbb{E}_{c\sim\mathcal{D}_{c},x_{1}\sim p_{\theta}(x_{1}\mid c)}[r(c,x_{1})]-\beta\mathbb{D}_{\mathrm{KL}}[p_{\theta}(x_{1}\mid c)\parallel p_{\mathrm{ref}}(x_{1}\mid c)]$ (2)
where the hyperparameter $\beta$ controls KL-regularization strength. RLHF training requires the sampled $x_{1}$ be differentiable. As shown in the Equation (3), the sampling result $x_{1}$ of flow matching can be written as the sum of the outputs of the model at each step.
$x_{1}=x_{0}+\frac{1}{N}\sum_{i=0}^{N-1}v_{\theta}(x_{\frac{i}{N}},\frac{i}{N})$ (3)
Inspired by Wu et al. [50], we enable the gradient of the model to make $x_{1}$ differentiable and maximize the reward through backpropagation. Instead of enabling the gradient at equal intervals, we randomly sample $k$ steps carrying gradient from the timesteps of inference as shown in the Figure 8. Considering that the model for video generation is large and the sequence is long while the memory of the NPUs is limited, we ignore the KL constraint for the sake of computational efficiency.
Figure 8: Illustration of diffusion RLHF. An iteration in the flow matching sampling can be written as $x_{t_{k+1}}=x_{t_{k}}+v_{\theta}(x_{t_{k}},t_{k})$. High-order differentiation can be avoided by stopping the gradient on the input of the model. For the random $k$ steps in the sampling process, the model gradient is enabled. At this time, $x_{1}$ contains the gradient and the parameters of these $k$ steps can be optimized through backpropagation.
@@ 7.2 Asynchronous Encode Server
Figure 9: Asynchronous Encoding Server Architecture. The feature extraction and model training processes are distributed across separate computing clusters. Data and feature transfers are handled asynchronously via network requests to optimize efficiency and resource utilization.
Since the text-to-video diffusion model performs denoising in the latent space, it requires both the VAE and text encoder to extract latent features and text embeddings, respectively. Although this inference stage involves no gradient computation, the T5-XXL [38] text encoder and VAE encoder still demand significant memory resources. In SD3.5L [12], feature extraction is performed offline, which inevitably consumes substantial storage space. To address these resource constraints, we propose a decoupled architecture that separates computing resources for inference and training, as shown in the Figure 9. Specifically, the VAE and text encoder operate on dedicated inference clusters, continuously generating processed input data (including image and video latent features and corresponding text embeddings) for DiT training. The training cluster then retrieves these pre-computed features via network transmission and focuses exclusively on the training process.